# ExLM: Rethinking the Impact of [MASK] Tokens in Masked Language Models

### v2

@@ Abstract
Masked Language Models (MLMs) have achieved remarkable success in many self-supervised representation learning tasks. MLMs are trained by randomly replacing some tokens in the input sentences with [MASK] tokens and predicting the original tokens based on the remaining context. This paper explores the impact of [MASK] tokens on MLMs. Analytical studies show that masking tokens can introduce the corrupted semantics problem, wherein the corrupted context may convey multiple, ambiguous meanings. This problem is also a key factor affecting the performance of MLMs on downstream tasks. Based on these findings, we propose a novel enhanced-context MLM, ExLM. Our approach expands [MASK] tokens in the input context and models the dependencies between these expanded states. This expansion increases context capacity and enables the model to capture richer semantic information, effectively mitigating the corrupted semantics problem during pre-training. Experimental results demonstrate that ExLM achieves significant performance improvements in both text modeling and SMILES modeling tasks. Further analysis confirms that ExLM enhances semantic representations through context enhancement, and effectively reduces the multimodality problem commonly observed in MLMs.
@@ 1 Introduction
Pre-trained language models based on the Masked Language Model (MLM) training objective have achieved significant success across various types of sequence data, including text (Devlin, 2018; Liu, 2019; Lan, 2019; He et al., 2020; Joshi et al., 2020; Meng et al., 2023), small molecules(Wang et al., 2019; Chithrananda et al., 2020; Pan, 2023; Zheng et al., 2024a), and proteins (Elnaggar et al., 2021; Lin et al., 2022, 2023; Zheng et al., 2024b), establishing themselves as a foundational approach for sequence representation learning. To enable the effective extraction of useful semantic information, MLM employs a mask-then-predict training strategy. During training, a certain proportion (typically $15\%$) of tokens in the input sequence are randomly replaced with a special symbol, [MASK]. The model is then trained to predict the original tokens based on the corrupted context. This process allows MLMs to learn meaningful semantic information in a self-supervised manner from large-scale unlabeled data. The learned knowledge can be effectively transferred to a wide range of downstream tasks, such as text classification and molecular property prediction, significantly improving the performance of deep learning models in these domains.
Figure 1: Illustrations of the vanilla MLM (a) and ExLM (b). MLM can be affected by the multimodality problem. In ExLM, the model creates multiple hidden states for each [MASK] token (e.g., $[\texttt{M}_{1,1}],[\texttt{M}_{1,2}],[\texttt{M}_{2,1}],[\texttt{M}_{2,2}]$). By leveraging a larger semantic space and explicitly modeling the dependencies between these states, the model can capture richer semantic information while mitigating the effects of multimodality.
Although MLMs have achieved significant success in numerous tasks and applications, their effectiveness remains an important research problem. During the MLM training process, parts of the input are masked, introducing unreal tokens ([MASK]) into the context. This process impacts MLMs from two critical aspects:
@@
These two aspects of impact are closely tied to the mask ratio used during pre-training. Higher mask ratios exacerbate the problems of unreal tokens and corrupted semantics at the same time, leading to a noticeable performance drop when the mask ratio is too high. While previous studies have investigated the unreal token issue and its impact on MLM performance (Clark, 2020; Meng et al., 2023), they have largely overlooked the corrupted semantics problem. Consequently, there has little work systematically exploring the impact of both issues on MLM performance, or comparing the extent of their effects separately. This gap arises because these two factors are interdependent and both rely on the mask ratio, making it challenging to design experiments that disentangle their respective effects.
To address this research gap, we conducted carefully-designed analytical experiments to separately evaluate the relative impact of these two factors. The experimental results demonstrate that the corrupted semantics problem has a significantly greater impact on MLM performance than the unreal tokens problem. This is further reflected in the increased multimodality problem (Gu et al., 2017) that arises when the model predicts the original tokens. Based on these findings, we propose a novel pre-training model, ExLM, which enhances context representation in MLMs. By expanding each [MASK] token in the input context into multiple hidden states, the model is provided with a larger semantic space. This allows it to capture richer semantic information associated with each [MASK] token, thereby mitigating the multimodality problem in token prediction. Furthermore, we introduce a transition matrix between these expanded states, enabling the model to directly capture semantic relationships among different states. To efficiently train the ExLM, we propose a state alignment algorithm based on dynamic programming, which aligns target tokens with expanded states in a data-driven manner, significantly improving model performance.
@@
  • We conduct the first systematic analysis of MLM behavior from the dual perspectives of both unreal tokens and corrupted semantics. Through a series of carefully designed experiments, we reveal that the corrupted semantics issue has a greater impact on MLM performance, providing new insights into MLM studies.
  • Based on the analysis, we propose ExLM, a novel pre-training model that enhances context representation for MLMs. By expanding [MASK] tokens into multiple states and employing a state alignment algorithm, ExLM effectively improves semantic modeling and mitigates the multimodality problem.
  • Extensive experimental results on text and molecular property prediction tasks demonstrate the superior performance of ExLM. Further analysis confirms its effectiveness and highlight its potential for addressing challenges in MLM pre-training.
Figure 2: Illustrations of the Repeated MLM experiment. Each token in the model input is repeated before being masked. The artificial redundancy introduced into the input ensures that replacing a token with [MASK] does not necessarily lead to semantic corruption in the context, as the repeated tokens provide additional information to preserve the original semantics.
@@ 2 Background of MLM
The pre-training objective of MLMs is to predict missing tokens from a partially masked input sequence. In this process, a portion of the input tokens is randomly selected and masked (i.e., replaced with [MASK]), forming a corrupted sequence. The model’s task is to recover the original tokens based on the remaining unmasked context as input.
Let’s define a sequence of tokens as $\mathbf{X}=[x_{1},x_{2},\dots,x_{n}]$. In the MLM pre-training process, some of the tokens in $\mathbf{X}$ are randomly replaced with a special token [MASK], producing a partially masked input sequence $\tilde{\mathbf{X}}=[\tilde{x}_{1},\tilde{x}_{2},\dots,\tilde{x}_{n}]$, where:
@@
3 Understanding the Impact of [MASK]
In this section, we design analytical experiments to explore the impact of unreal tokens and corrupted semantics on MLM performance. However, since these two factors are strongly interrelated in MLMs, we first designed a Repeated MLM experiment to decouple unreal tokens and corrupted semantics (Section 3.1). This allows us to separately explore the impact of each factor and demonstrate that corrupted semantics has a significantly greater effect on MLM performance (Section 3.2). Additionally, our analysis further shows that more severe corrupted semantics lead to the loss of critical context semantics, which exacerbates the multimodality problem in token predictions (Section 3.3).
3.1 Decoupling Corrupted Semantics and Unreal Tokens
@@
The core idea of this experiment is to artificially introduce redundancy into the MLM’s input, ensuring that replacing a token with [MASK] does not necessarily result in semantic corruption in the context.
Specifically, as illustrated in Figure 2, before feeding the sequence into the MLM, we first repeat each token $k$ times (where $k$ is a hyperparameter set based on experimental requirements) and then randomly mask the repeated sequence at a certain ratio $p$, selecting $p$ of the tokens to replace with [MASK]. This creates an interesting phenomenon: while the proportion of unreal [MASK] tokens in the context remains $p$ due to the masking ratio, the proportion of corrupted semantics changes. Since each token has $k$ copies and the probability of each copy being masked is $p$, the probability of the semantic information carried by a token being completely corrupted becomes $p^{k}$.
By fixing $p$ and varying $k$, we can ensure that the proportion of [MASK] tokens in the context remains constant while the degree of corrupted semantics varies. This enables us to control variables and systematically compare the impact of these two factors on MLM performance separately.
@@
3.2 What Matters More: Corrupted Semantics or Unreal Tokens
Under the experimental setup designed in Section 3.1, we have trained a series of MLMs with different repetition times $k$ and mask ratios $p$ to analyze MLM’s behavior. For consistency, all training hyperparameters of the MLMs in the experiments are kept the same except for $p$ and $k$. Additionally, during downstream fine-tuning, the input is repeated with the same repetition times $k$ as in pre-training. More detailed training configurations and hyperparameters can be found in Appendix B. The results of the MNLI task, evaluated using accuracy as the primary metric (Williams et al., 2018), are presented in Figure 3. We also provide the results of this experiments on more tasks in Appendix F. Some entries are blank (gray areas) due to overly low proportions of corrupted semantics (less than 0.05%), causing very low loss during pre-training and unstable model training. For comparison, the proportions of corrupted semantics for each experiment are provided in Figure 4.
@@ 3.3 Core Impact of Corrupted Semantics: Multimodality
We conducted a deeper analysis to investigate how Corrupted Semantics influence the performance of MLMs. As illustrated in Figure 1(a), the reconstruction of a [MASK] token relies on the semantic information provided by its context. When a corrupted context is given, it may imply multiple different potential semantics, resulting in significantly different reconstruction outcomes. This phenomenon is referred to as multimodality, a concept initially proposed in NAT (Gu et al., 2017). Multimodality causes MLMs to produce more mixed and uncertain predictions during pretraining, thereby significantly affecting downstream performance. To gain a more straightforward understanding of multimodality, we analyzed the prediction entropy of MLMs with different repetition times $k$ and mask ratios $p$. We also have provided more details on the entropy calculation process in Appendix E. As illustrated in Figure 5, with an increase in semantic corruption, the prediction entropy also rises, indicating a more severe multimodality problem. Such multimodality problem also significantly impacts the performance of MLMs.
On the other hand, when corruption is too low (e.g., 2.25%), the context still contains plenty of semantic clues, making token prediction too easy. At this point, the model’s predictions also exhibit very low uncertainty. The training curve analysis in Appendix G also shows that when the corruption is too low, the training task becomes overly simplistic. This simplicity prevents the model from learning deeper knowledge and ultimately reduces performance.
Therefore, an optimal approach would be to design the input context with a reasonable degree of semantic corruption while ensuring the model can correctly handle the missing semantics. This balance would avoid the negative impacts of the semantic multimodality problem while maintaining the model’s performance.
@@ 4.2 Modeling Semantics with Enhanced Context
This expanded sequence $\mathbf{X^{\prime}}$, consisting of embeddings for both the original tokens and the duplicated [MASK] tokens, is then passed into a Transformer Encoder $\boldsymbol{\theta}$ for contextual encoding (Vaswani et al., 2017):
$\mathbf{H^{\prime}}=[\mathbf{h}_{x_{1}},\mathbf{h}_{x_{2}},\dots,\mathbf{h}_{[{\rm MASK}]}^{(1)},\dots,\mathbf{h}_{[{\rm MASK}]}^{(k)},\dots,\mathbf{h}_{x_{n}}],$
@@
$\mathbf{E}=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d}}+\mathbf{M}\right),$
@@
$\mathbf{P}=\text{softmax}(\mathbf{H}\mathbf{W}_{\text{P}}^{T}),$
@@ 4.3 Pre-training of ExLM: States Alignment
Formulating States Alignment as a DAG decoding problem.
Our goal is to maximize the probability of the DAG decoding all target tokens $\mathbf{Y}$. Formally, let $\Gamma$ denote the set of all possible paths (i.e., all possible ways to align states to the target tokens). Each path $\mathbf{A}\in\Gamma$ represents one particular alignment of states to the target tokens in $\mathbf{Y}$. The training objective can thus be written as a marginalization over all possible alignments:
where $\mathbf{X^{\prime}}$ is the input sequence (with enhanced context), $\mathbf{\theta}$ denotes the model parameters, and $P_{\theta}(\mathbf{Y},\mathbf{A}\mid\mathbf{X^{\prime}})$ is the probability of generating the target sequence $\mathbf{Y}$ via a particular alignment $\mathbf{A}$.
This objective requires summing over all possible alignment paths $\mathbf{A}\in\Gamma$, which can be computationally expensive. To address this, we employ a dynamic programming (DP) algorithm that efficiently computes this sum with a complexity of $\mathcal{O}(M)$, where $M$ is the length of the target sequence. This approach significantly improves computational efficiency. Details of the DP algorithm are provided in Appendix H and the efficiency analysis is provided in Appendix I.

### v3

@@ Abstract
Masked Language Models (MLMs) have achieved remarkable success in many self-supervised representation learning tasks. MLMs are trained by randomly masking portions of the input sequences with [MASK] tokens and learning to reconstruct the original content based on the remaining context. This paper explores the impact of [MASK] tokens on MLMs. Analytical studies show that masking tokens can introduce the corrupted semantics problem, wherein the corrupted context may convey multiple, ambiguous meanings. This problem is also a key factor affecting the performance of MLMs on downstream tasks. Based on these findings, we propose a novel enhanced-context MLM, ExLM. Our approach expands [MASK] tokens in the input context and models the dependencies between these expanded states. This enhancement increases context capacity and enables the model to capture richer semantic information, effectively mitigating the corrupted semantics problem during pre-training. Experimental results demonstrate that ExLM achieves significant performance improvements in both text modeling and SMILES modeling tasks. Further analysis confirms that ExLM enriches semantic representations through context enhancement, and effectively reduces the semantic multimodality commonly observed in MLMs.
@@ 1 Introduction
Pre-trained masked language models (MLMs) have achieved significant success across various types of sequence data, including text (Devlin, 2018; Liu, 2019; Lan, 2019; He et al., 2020; Joshi et al., 2020; Meng et al., 2023), small molecules (Wang et al., 2019; Ross et al., 2022; Pan, 2023; Zheng et al., 2024a), and proteins (Lin et al., 2022, 2023; Su et al., 2023; Zheng et al., 2024b; Hayes et al., 2025), establishing themselves as a foundational approach for sequence representation learning tasks. To enable the effective extraction of useful semantic information, MLM employs a mask-then-predict training strategy. During the pre-training process, a certain proportion (typically $15\%$) of tokens in the input sequence are randomly replaced with a special symbol, [MASK]. The model is then trained to predict the original tokens based on the corrupted context. This process allows MLMs to learn meaningful semantic information in a self-supervised manner from large-scale unlabeled data. The learned knowledge can be effectively transferred to a wide range of downstream tasks, such as text classification and molecular property prediction, significantly improving the performance of deep learning models in these domains.
Figure 1: Illustrations of the vanilla MLM (a) and ExLM (b). MLM can be affected by the multimodality problem. In ExLM, the model creates multiple hidden states for each [MASK] token (e.g., $[\texttt{M}_{1,1}],[\texttt{M}_{1,2}],[\texttt{M}_{2,1}],[\texttt{M}_{2,2}]$). By leveraging a larger semantic space and explicitly modeling the dependencies between these states, the model can capture richer semantic information in the enhanced context while mitigating the effects of multimodality.
Although MLMs have achieved significant success in numerous tasks and applications, their effectiveness remains an important research problem. During the MLM training process, parts of the input are masked, and unreal tokens ([MASK]) are introduced into the context. This process impacts the input context of MLMs in two critical aspects:
@@
These two aspects of impact are closely tied to the mask ratio used during pre-training. Higher mask ratios exacerbate the problems of unreal tokens and corrupted semantics at the same time, leading to a noticeable performance drop when the mask ratio is too high. While previous studies have investigated the unreal token problem and its impact on MLM performance (Clark, 2020; Meng et al., 2023), they have largely overlooked the corrupted semantics problem. Consequently, there is little work systematically exploring the impact of both problems on MLM performance, or evaluating the magnitude of their effect independently. This gap arises because these two factors are interdependent and both rely on the mask ratio, making it challenging to design experiments that disentangle their respective effects.
To fill this research gap, we designed and conducted the Repeated MLM analytical experiments (shown in Figure 2) to separately evaluate the relative impact of these two factors. The experimental results demonstrate that the corrupted semantics problem has a significantly greater impact on MLM performance than the unreal tokens problem. This is further reflected in a stronger semantic multimodality (Gu et al., 2017), where multiple plausible predictions exist for the original tokens due to the ambiguous or context-dependent meanings. Based on these findings, we propose a novel pre-trained model, ExLM, which enhances context representation in MLMs. By expanding each [MASK] token in the input context into multiple hidden states, the model is provided with a larger semantic space. This allows it to capture richer semantic information associated with each [MASK] token, thereby reducing the semantic multimodality in token prediction. Furthermore, we introduce a transition matrix between these expanded states, enabling the model to directly capture semantic relationships among different states. To efficiently train the ExLM, we propose a state alignment algorithm based on dynamic programming, which aligns target tokens with expanded states in a data-driven manner, significantly improving model performance.
@@
  • We conduct the first systematic analysis of MLM behavior from the dual perspectives of both unreal tokens and corrupted semantics. Through a series of carefully designed experiments, we reveal that the corrupted semantics problem has a greater impact on MLM performance, providing new insights into MLM studies.
  • Based on the analysis, we propose ExLM, a novel pre-trained model that enhances context representation for MLMs. By expanding [MASK] tokens into multiple states and employing a state alignment algorithm, ExLM effectively improves semantic modeling and reducing the semantic multimodality in the context.
  • Extensive experimental results on text and molecular property prediction tasks demonstrate the superior performance of ExLM. Further analysis confirms its effectiveness and highlights its potential for addressing challenges in MLM pre-training.
[…]
@@ 2 Background of MLM
The pre-training objective of MLMs is to predict missing tokens from a partially masked input sequence. In this process, a portion of the input tokens is randomly selected and masked (i.e., replaced with [MASK]), forming a corrupted sequence. The model’s task is to recover the original tokens using the remaining unmasked context as input.
Specifically, let’s define a sequence of tokens as $\mathbf{X}=[x_{1},x_{2},\dots,x_{n}]$. In the MLM pre-training process, some of the tokens in $\mathbf{X}$ are randomly replaced with a special token [MASK], producing a partially masked input sequence $\tilde{\mathbf{X}}=[\tilde{x}_{1},\tilde{x}_{2},\dots,\tilde{x}_{n}]$, where:
@@
3 Understanding the Impact of [MASK]
In this section, we design analytical experiments to explore the impact of unreal tokens and corrupted semantics on MLM performance. However, since these two factors are strongly interrelated in MLMs, we first designed a Repeated MLM experiment to decouple unreal tokens and corrupted semantics (Section 3.1). This allows us to separately explore the impact of each factor and demonstrate that corrupted semantics has a significantly greater effect on MLM performance (Section 3.2). Additionally, our analysis further shows that more severe corrupted semantics lead to the loss of critical context semantics, which exacerbates the multimodality in token predictions (Section 3.3).
3.1 Decoupling Corrupted Semantics and Unreal Tokens
@@
The core idea of this experiment is to artificially introduce redundancy into the MLM’s input, ensuring that replacing a token with [MASK] does not inevitably lead to semantic corruption in the context.
Specifically, as illustrated in Figure 2, before feeding the sequence into the MLM, we first repeat each token $k$ times, where $k\in\mathbb{N}$ is a hyperparameter set based on experimental requirements. We then randomly mask the repeated sequence at a certain ratio $p$, which means replacing $p\in(0.0\%,100.0\%)$ of the tokens with [MASK]. This setup creates an interesting phenomenon: while the proportion of unreal [MASK] tokens in the context remains $p$ due to the masking ratio, the proportion of corrupted semantics changes. Since each token has $k$ copies and the probability of each copy being masked is $p$, the probability of the semantic information carried by a token being completely corrupted becomes $p^{k}$. We also have provided a detailed proof about the results in Appendix C.
By keeping $p$ fixed and varying $k$, we ensure that the proportion of [MASK] tokens in the context remains constant while adjusting the degree of corrupted semantics. This enables us to control variables effectively, and systematically compare the impact of these two factors on MLM performance by measuring each factor separately.
3.2 What Matters More: Corrupted Semantics or Unreal Tokens
@@
Under the experimental setup designed in Section 3.1, we have trained a series of MLMs with different repetition times $k$ and mask ratios $p$ to analyze MLM’s behavior. For consistency, all training hyperparameters of the MLMs in the experiments are kept the same except for $p$ and $k$. Additionally, during downstream fine-tuning, the input is repeated with the same repetition times $k$ as in pre-training. More detailed training configurations and hyperparameters can be found in Appendix B. The results of the MNLI task, evaluated using accuracy as the primary metric (Williams et al., 2018), are presented in Figure 3. We also provide the results of this experiments on more tasks in Appendix F. Some entries are blank (gray areas) due to overly low proportions of corrupted semantics (less than $0.05\%$), causing very low loss during pre-training and unstable model training problem. For comparison, the proportions of corrupted semantics for each experiment are provided in Figure 4.
@@ 3.3 Core Impact of Corrupted Semantics: Multimodality
We conducted a deeper analysis to investigate how corrupted semantics influence the performance of MLMs. As illustrated in Figure 1(a), the reconstruction of a [MASK] token relies on the semantic information provided by its context. When a corrupted context is given, it may imply multiple different potential semantics, resulting in significantly different reconstruction outcomes. This phenomenon is referred to as multimodality, a concept initially proposed in NAT (Gu et al., 2017). Multimodality causes MLMs to produce more mixed and uncertain predictions during pre-training, thereby significantly affecting downstream performance. To gain a more straightforward understanding of multimodality, we analyzed the prediction entropy of MLMs with different repetition times $k$ and mask ratios $p$. We also have provided more details on the entropy calculation process in Appendix E. As illustrated in Figure 5, with an increase in semantic corruption, the prediction entropy also rises, indicating a more severe semantic multimodality. Such multimodal phenomenon also significantly impacts the performance of MLMs.
On the other hand, when corruption is too low (e.g., $2.25\%$), the context still contains plenty of semantic clues, making token prediction too easy. At this point, the model’s predictions also exhibit very low uncertainty. The training curve analysis in Appendix G also shows that when the corruption is too low, the training task becomes overly simplistic. This simplicity prevents the model from learning deeper knowledge and ultimately reduces performance.
Therefore, an optimal approach would be to design the input context with a reasonable degree of semantic corruption, ensuring the model can correctly handle the missing semantics and learn meaningful knowledge. This balance would avoid the negative impacts of the semantic multimodality problem while maintaining the model’s performance.
@@ 4.2 Modeling Semantics with Enhanced Context
This expanded sequence $\mathbf{X^{\prime}}$, consisting of embeddings for both the original tokens and the duplicated [MASK] tokens, is then passed into a Transformer Encoder $\boldsymbol{\theta}$ for contextual encoding $\mathbf{H}$ (Vaswani et al., 2017):
$\mathbf{H}=[\mathbf{h}_{x_{1}},\mathbf{h}_{x_{2}},\dots,\mathbf{h}_{[{\rm MASK}]}^{(1)},\dots,\mathbf{h}_{[{\rm MASK}]}^{(k)},\dots,\mathbf{h}_{x_{n}}],$
@@
$\mathbf{E}=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}+\mathbf{M}\right),$
@@
$\mathbf{P}=\text{softmax}(\mathbf{H}\mathbf{W}_{\text{P}}^{\top}),$
@@ 4.3 Pre-training of ExLM: States Alignment
Formulating States Alignment as a DAG decoding problem.
Our goal is to maximize the probability of the DAG decoding all target tokens $\mathbf{Y}$ (i.e., all masked tokens). Formally, let $\Gamma$ denote the set of all possible paths (i.e., all possible ways to align states to the target tokens). Each path $\mathbf{A}\in\Gamma$ represents one particular alignment of states to the target tokens in $\mathbf{Y}$. The training objective can thus be written as a marginalization over all possible alignments:
where $\mathbf{X^{\prime}}$ is the input sequence (with enhanced context), $\mathbf{\theta}$ denotes the model parameters, and $P_{\theta}(\mathbf{Y},\mathbf{A}\mid\mathbf{X^{\prime}})$ represents the probability of generating the target sequence $\mathbf{Y}$ through a specific alignment $\mathbf{A}$.
This objective requires summing over all possible alignment paths $\mathbf{A}\in\Gamma$, which can be computationally expensive. To address this, we employ a dynamic programming (DP) algorithm that efficiently computes this sum with a complexity of $\mathcal{O}(M)$, where $M$ is the number of all masked tokens. This approach significantly improves computational efficiency. Details of the DP algorithm are provided in Appendix H and the efficiency analysis is provided in Appendix I.
5 Experimental Results
We evaluate the performance of the ExLM model in both text modeling and SMILES modeling tasks. SMILES is a sequential representation of molecular information, and a more detailed explanation can be found in Appendix D.
We pre-train ExLM on data of text and SMILES separately. Then we fine-tune and evaluate ExLM across diverse benchmarks and verify the contribution of each component through ablation studies. Finally, a visualization analysis is included to explain the advantages of unified modeling.
@@ 5 Experimental Results
5.1 Results on SMILES Representation Learning
Pre-training.
Fine-tuning.
Results.
5.2 Results on Textual Representation Learning
Pre-training.
For textual pre-training, we adopt the English Wikipedia and BookCorpus datasets (Devlin, 2018) as the pre-training dataset. The model size of ExLM is consistent with BERT base (Devlin, 2018). We also train ExLMLARGE with the same size as BERT large. For more details about pre-training settings, please see Appendix N.
Fine-tuning.
Results.
As shown in Table 3, we evaluate ExLM on the GLUE and SQuAD 2.0 development sets. ExLM achieves the best performance on SQuAD 2.0 and 7 GLUE tasks, and also performs very closely to the best baseline model on the remaining one GLUE task. Furthermore, ExLM demonstrates significantly higher average performance compared to other baseline models. Notably, RoBERTa in the table uses the same pre-training data and settings as ExLM, yet ExLM shows significant improvements. This demonstrates the superior performance and effectiveness of ExLM. ExLMLARGE also achieves obvious performance improvements compared to ExLM due to the increased scale.
[…]
ExLM w/o 2D RoPE 84.6 91.1 91.3 56.7 80.9
ExLM w/o Transitions 83.8 90.9 91.1 55.6 80.4
Vanilla MLM 83.6 90.0 90.3 54.7 79.6
Vanilla MLM++ 84.4 91.2 90.6 56.3 80.7
ExLM 85.1 91.4 91.3 57.6 81.4